OpenStudy (anonymous):

Solve: | m + 8| = 16 Answer {8} {-24, 8} {8, 16} {-24}

jimthompson5910 (jim_thompson5910):

| m + 8| = 16 m + 8 = 16 or m + 8 = -16 Now solve each equation for 'm'